OpenStudy (anonymous):

What is the solution to the equation? (x-16)+5=-15 x=?

OpenStudy (midhun.madhu1987):

x - 16 + 5 = -15 x - 11 = -15 Now add 11 to both sides of the equation.. what u get?

OpenStudy (anonymous):

You got this @LOLAri

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

-4

OpenStudy (phi):

almost you should get x-11 + 11 = -15 + 11

OpenStudy (anonymous):

You got it just for got to place to dash

OpenStudy (phi):

for -15+11 you can do 15-11 and then make the answer negative.

OpenStudy (anonymous):

forgot*

OpenStudy (anonymous):

its ok